a sample of impure zinc dust of mass 0.350g was added tO an excess of aqueous iro(11) sulphate. the iron produced required 100cubic centimeters of 0.0200decimeter cubed potassium manganate(7) for titration.
write the equation for the reactions take which take place?
calculate the percentage purity of zinc dust?
1
Expert's answer
2020-06-02T14:07:34-0400

1) zinc dust: Zn + FeSO4 = ZnSO4 + Fe

2) titration: Fe + KMnO4 + 2 H2O = Fe(OH)3 + MnO2 + KOH

calculation:

number of moles of KMnO4 = 0.1 L (i.e. 100 cubic centimeters) * 0.0200 M (i.e. moles per decimeter cubic) = 0.0020 moles = number of moles of Fe = number of moles of Zn

mass of Zn = 0.0020 mole * 65.38 g/mole = 0.131 g

percentage purity if 0.131 g / 0.350 g * 100% = 37.4 %